# Batch speculative decoding Done right

### v1

@@ Abstract
Speculative decoding speeds up LLM inference by using a small draft model to propose multiple tokens that a target model verifies in parallel. Extending this idea to batches is essential for production serving, but it introduces the ragged tensor problem: sequences in the same batch accept different numbers of draft tokens, breaking right-alignment and corrupting position IDs, attention masks, and KV-cache state. We show that several existing batch implementations violate output equivalence—the fundamental requirement that speculative decoding must produce identical token sequences to standard autoregressive generation. These violations occur precisely due to improper handling of the ragged tensor problem. In response, we (1) characterize the synchronization requirements that guarantee correctness, (2) present a correctness-first batch speculative decoding EqSpec that exposes realignment as consuming 40% of overhead, and (3) introduce EXSpec, which maintains a sliding pool of sequences and dynamically forms same-length groups, to reduce the realignment overhead while preserving per-sequence speculative speedups. On the SpecBench dataset, across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B target/draft pairs, our approach achieves up to 3× throughput improvement at batch size 8 compared to batch size 1, with efficient scaling through batch size 8, while maintaining 95% output equivalence. Our method requires no custom kernels and integrates cleanly with existing inference stacks. Our code is available at https://github.com/eBay/spec_dec.
@@ 1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) accelerates LLM inference by using a small draft model to propose multiple tokens that the target model verifies in parallel, shifting from memory-bound sequential generation to compute-intensive verification and delivering single-sequence speedups (Xia et al., 2024). Batch speculative decoding aims to combine this per-sequence acceleration with standard batching by processing multiple sequences (batch dimension) while verifying multiple draft tokens per sequence (sequence dimension). The core challenge is the ragged-tensor effect (Qian et al., 2024): in each verification round, sequences accept a single series of the proposed draft tokens, but prefix lengths differ across sequences (e.g., one accepts five tokens while another accepts one), misaligning sequence lengths. This raggedness violates the rectangular-tensor assumption required for GPU-parallel execution.
Figure 1 highlights a critical, often overlooked issue in batch speculative decoding: methods with impressive throughput can produce corrupted outputs. For lossless acceleration, speculative decoding must yield identical outputs to standard autoregressive generation (Leviathan et al., 2023). As a reference point, the widely used HuggingFace implementation (Wolf et al., 2020) preserves this guarantee, but only for batch size 1. By contrast, in our tests of public batch implementations—specifically, BSP (Su et al., 2023) and DSD (Yan et al., 2025)—this requirement is violated at batch sizes >1>1, manifesting as repetitive tokens or <unk> symbols under greedy decoding rather than matching standard generation.
@@
Figure 1: Batch speculative decoding on Vicuna-7B/68M: Existing methods achieve high throughput but violate the fundamental requirement of output equivalence by producing corrupted outputs. Our approach maintains perfect correctness while still achieving competitive performance.
These failures share a single root cause—broken synchronization invariants (position tracking, attention, KV-cache) across ragged tensors. We first propose EqSpec by formalizing these invariants and enforcing them with synchronization-aware scheduling (Section 3.1). Concretely, EqSpec specifies the minimal synchronization needed for correctness and shows that realignment accounts for 40% of computation—an inherent cost of maintaining invariants across ragged tensors. This fundamental overhead helps explain why even production systems struggle: vLLM’s speculative decoding underperforms its non-speculative baseline at higher batch sizes (leading to deprecation in its v1 engine), while SGLang with EAGLE (Li et al., 2024a) consistently exhibits negative speedups. Our analysis indicates that the superlinear growth of synchronization overhead is an inevitable cost of correctness in batch speculative decoding—a barrier no existing system has overcome.
@@
To address this in practice, our main algorithm EXSpec (Section 3.2) expands the scheduling scope: it maintains a sliding window of active sequences and dynamically groups those with identical lengths, eliminating realignment for homogeneous groups. This strategy preserves the scaling efficiency of standard batching—where GPU parallelism drives throughput—while retaining the per-sequence acceleration benefits of speculation. At batch size 8, we achieve a $3\times$ throughput improvement over batch size 1. Nevertheless, beyond batch size 8, throughput degrades as grouping success rates decline, forcing more frequent fallbacks to expensive realignment. Section 4 examines these scaling dynamics and the relationship between sequence diversity, grouping effectiveness, and alignment overhead.
@@
Our experiments on SpecBench (Xia et al., 2024) yield two main results. First, Correctness: unlike prior approaches such as BSP and DSD, which suffer severe output corruption, our method preserves approximately 95% output equivalence across Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024) model pairs. Second, Scalability: at batch size 8, EXSpec achieves up to a $3\times$ speedup over batch size 1. Our contributions are summarized as follows:
  • We provide a correctness-first analysis of the ragged-tensor problem in batch speculative decoding, identifying precise synchronization requirements for correctness and explaining why existing methods fail (Section 2).
  • We present a unified solution that maintains correctness through precise synchronization invariants while avoiding their overhead via cross-batch scheduling of same-length sequence groups (Section 3).
  • We experimentally demonstrate that our approach achieves both >>95% output correctness and positive scaling through batch size 8, successfully multiplying batch parallelism with per-sequence speculation gains, whereas production systems (vLLM, SGLang) exhibit negative speedups (Section 4).
@@ 2 Design Space Analysis
When sequences within a batch accept different numbers of draft tokens during verification, tensors become irregular, violating GPUs’ requirement for rectangular layouts—this is the ragged-tensor problem illustrated in Figure 2. Despite batching’s centrality to production deployments, existing implementations lack a principled design that preserves output equivalence with standard decoding while scaling with batch size. We identify three approaches to handle raggedness: Masking, Rollback, and Dynamic Padding. Yet, as our systematic analysis shows, current instantiations of these approaches fail to simultaneously maintain correctness and performance at scale. To close this gap, we first analyze the pitfalls of each approach and then introduce a correctness-first algorithmic design with explicit synchronization requirements for reliable batch speculative decoding.
@@
This approach operates directly on ragged tensors by masking rejected tokens in attention and reassigning position IDs so new tokens align with their content positions. Across verification rounds with varying rejections, sequences accumulate padding in various positions (middle and right), forming non-contiguous position IDs that standard Transformer implementations handle poorly. BSP (Su et al., 2023) attempts this via masking but fails to maintain position-ID consistency across iterations, yielding corrupted outputs (Figure 1). EAGLE’s experimental batching code [^1] While EAGLE’s main contribution concerns improved draft models rather than batching, its repository includes experimental batch-related code we analyzed for implementation challenges. https://github.com/SafeAILab/EAGLE/issues/250 (Li et al., 2025) encounters similar framework limitations. Supporting non-contiguous position IDs would require custom CUDA kernels for each base model (Qian et al., 2024)—a prohibitive engineering cost that sacrifices portability.
@@
After each verification step, all sequences are truncated to the batch’s minimum accepted length (Wolf et al., 2020; kamilakesbi, 2024). This guarantees alignment but discards correctly verified tokens from faster sequences. As batch size grows and acceptance-rate variance widens, the waste compounds; in the extreme, one persistently rejecting sequence forces single-token progress for the entire batch. In effect, throughput collapses to that of the worst-performing sequence, undermining speculative gains and rendering the approach impractical at larger scales.
Algorithm 1 BatchVerify: Single Forward Pass
1: Target model $\mathcal{M}_{t}$, Sequences $\mathcal{S}$, Draft tokens $D$, KV cache
2: Accepted tokens AA, Bonus tokens $B$
3: if first iteration then
4:   $\mathcal{X}\leftarrow\mathcal{S}\oplus D$
5: else
6:   $\mathcal{X}\leftarrow D$
7: $\textit{logits},\textit{KVCache}\leftarrow\mathcal{M}_{t}(\mathcal{X},\textit{KVCache})$
8: $\textit{pred\_tokens}\leftarrow\arg\max(\textit{logits, dim=vocab})$
9: $\triangleright$ Vectorized first mismatch detection
10: $\textit{matches}\leftarrow(\textit{pred\_tokens}=D)$
11: $J\leftarrow\arg\max(\neg\textit{matches, dim=seq})$
12: $\triangleright$ Ragged shape acceptance, no vectorization
13: for each sequence $i$ in batch do
14:   $A[i]\leftarrow D[i][:j]$
15:   $\triangleright$ Get bonus token from first mismatch
16:   $\textit{bonus\_logit}\leftarrow\textit{logits}[i,|\mathcal{S}[i]|+j]$
17:   $B[i]\leftarrow\arg\max(\textit{bonus\_logit})$
18: return $A,B,KVCache$
Algorithm 2 EqSpec
1: Draft model $\mathcal{M}_{d}$, Target model $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Max tokens $T$, Draft length $K$
2: Generated sequences $\mathcal{S}$
@@
3: $\mathcal{S}\leftarrow\texttt{Tokenize}(\mathcal{P})$ $\triangleright$ Batch left padding
4: $\textit{KVCache}\leftarrow\emptyset$
5: while until max new tokens do
6:   Phase 1: Draft Generation
7:   $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{S},K)$
8:   Phase 2: Batch Verification
9:   $A,B,\textit{KVCache}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{S},D,\textit{KVCache})$
10:   $\triangleright$ See Figure 3 for illustration on index offset.
11:   Phase 3: Unpad-Append-Repad
12:   for each sequence $i$ in batch do
13:    $\mathcal{S}[i]\leftarrow\texttt{Unpad}(\mathcal{S}[i])$
14:    $\mathcal{S}[i]\leftarrow\mathcal{S}[i]\oplus A[i]\oplus B[i]$
@@
17: return $\mathcal{S}$
✓ Dynamic Padding Approach.
This approach realigns sequences after each verification by adjusting left padding to maintain right alignment, preserving all accepted tokens. While conceptually simple, correctness requires tight synchronization of position IDs, attention masks, and the KV-cache. DSD’s experimental code (Yan et al., 2025) follows this idea but merely repads at each step—adding varying left padding without ever unpadding—thereby inflating sequences. It also contains three critical errors: (i) sampling bonus tokens from the draft model rather than the target model; (ii) redundantly regenerating KV-cache entries, causing memory bloat; and (iii) desynchronizing padding, position IDs, and the KV-cache across iterations. Despite the overhead of repeated realignment, a correct dynamic-padding implementation fits within standard frameworks and preserves all verified tokens.
Among the three approaches, only dynamic padding is viable: position-ID schemes require custom kernels that undermine portability, and rollback discards verified tokens at rates that grow with batch size; dynamic padding maintains correctness within standard frameworks.
3 Method
We present a synchronization-aware approach to batch speculative decoding that co-designs correctness and efficiency. The core tension is that preserving correctness requires synchronizing position IDs, attention masks, and the KV-cache across ragged tensors—an overhead that can consume 40% of computation. We introduce two complementary mechanisms: EqSpec specifies and enforces the minimal synchronization invariants required for valid computation (Section 3.1), while EXSpec groups same-length sequences to avoid synchronization overhead (Section 3.2). Together, these form a unified system in which correctness constraints drive scheduling, enabling both output equivalence and practical performance.
Figure 3: EqSpec synchronizes via unpad–append–repad.
3.1 Minimal Batch Realignment: EqSpec
Figure 3 illustrates the core challenge and our remedy for maintaining correctness in batch speculative decoding. After each verification round, sequences accept different numbers of draft tokens, producing ragged tensors that GPUs cannot process. To restore a valid batch, we apply an unpad–append–repad procedure that converts ragged outputs back to a rectangular layout while preserving three invariants: contiguous position IDs, valid attention masks, and aligned KV-cache entries.
Correct implementation requires padding-agnostic position IDs that reset to zero at the first content token, attention masks that exclude padding tokens, and precise KV-cache realignment after each verification step (Algorithm 1). After unpad–append–repad, index offsets shift across sequences; consequently, position IDs and attention masks must be recomputed to preserve correct token relationships. Moreover, the bonus token introduces a special case: it is sampled from the target model’s output distribution at the first mismatch position after the verification forward pass has already completed. Consequently, while it encodes the target’s authoritative correction, it lacks KV-cache entries in either the draft or target model because KV-cache is computed only for tokens that were present in the input sequence during the forward pass—and the bonus token, being an output, was not yet part of that input sequence. Therefore, it must be included in the next forward pass to create its KV-cache entries, further complicating synchronization. Finally, the realignment is resource-intensive: the KV-cache consists of rank-4 tensors $(\text{batch}\times\text{heads}\times\text{sequence}\times\text{dimension})$, and each padding adjustment triggers allocation and concatenation of high-dimensional zeros. Algorithm 2 details the complete EqSpec procedure.
@@ 3.1 Minimal Batch Realignment: EqSpec
Speedup Analysis.
The speedup of speculative decoding depends on both the token-acceptance rate and computational costs. The original formulation by Leviathan et al. (2023) analyzes single-sequence performance, modeling the expected tokens generated per iteration as $(1-\alpha^{k+1})/(1-\alpha)$, where $\alpha$ is the token acceptance rate (TAR) and $k$ is the number of draft tokens per speculation round. This single-sequence view assumes negligible batch overhead and focuses purely on acceptance dynamics.
Batch speculative decoding, however, introduces additional complexities not captured by that model—most notably alignment overhead, which can dominate and degrade performance. We therefore introduce a batch-aware speedup model:

### v2

@@ Abstract
Speculative decoding must produce outputs distribution identical to standard autoregressive generation—this output equivalence is not an optimization target but the defining criterion of valid speculative decoding. We demonstrate that all existing batch speculative decoding implementations violate this fundamental requirement, producing corrupted outputs ranging from repetitive tokens to gibberish. These failures stem from the ragged tensor problem: sequences in the same batch accept different numbers of draft tokens, desynchronizing position IDs, attention masks, and KV-cache state. We present the first authentic batch speculative decoding framework. We (1) formalize the synchronization invariants that valid batch speculative decoding must satisfy, (2) present EQSPEC, the first algorithm that guarantees output equivalence, and analyze its cost structure to show that alignment overhead grows superlinearly and consumes up to 40% of computation, and (3) introduce EXSPEC, which reduces this overhead through cross-batch scheduling that dynamically groups same-length sequences. On SpecBench across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B pairs, our methods achieve up to 3× throughput improvement at batch size 8 while maintaining algorithmic correctness. Our methods achieve 95% decoding-equivalence, with residual divergence attributable to floating-point non-determinism in GPU inference, not the synchronization failures that cause near-zero equivalence of prior methods. Our code is available at https://github.com/eBay/spec_dec
Large Language Models, Speculative Decoding, Inference Optimization, Batch Processing
@@ 1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) accelerates LLM inference by using a small draft model to propose multiple tokens that the target model verifies in parallel. The fundamental requirement of speculative decoding is output equivalence (Leviathan et al., 2023; Xia et al., 2024): the generated token distribution must remain identical to standard autoregressive generation. Any method that violates this requirement, regardless of its throughput, does not constitute valid speculative decoding. Batch speculative decoding aims to combine this per-sequence acceleration with standard batching, but as we demonstrate, all existing implementations fail to preserve output equivalence when batch size exceeds one.
Figure 1 highlights a fundamental but systematically overlooked failure in batch speculative decoding: methods with impressive throughput can produce corrupted outputs. The widely used HuggingFace implementation (Wolf et al., 2020) preserves output equivalence, but only for batch size 1. By contrast, in our tests of public batch implementations—specifically, BSP (Su et al., 2023) and DSD (Yan et al., 2025), produce repetitive tokens or <unk> symbols under greedy decoding rather than matching standard generation. These are not minor quality degradations but fundamental violations of decoding equivalence. Equivalence to reference decoding is a yes/no property: prior batch implementations fail to preserve the target model’s output distribution, yielding near-zero exact-match to standard autoregressive decoding due to synchronization errors. The root cause is improper handling of the ragged tensor problem (Qian et al., 2024), which desynchronizes position IDs, attention masks, and KV-cache state across sequences in a batch.
@@
Figure 1: Batch speculative decoding on Vicuna-7B/68M: Existing methods achieve high throughput but violate the fundamental requirement of output equivalence by producing corrupted outputs. Our approach preserves reference-decoding equivalence while still achieving competitive performance.
Valid batch speculative decoding requires maintaining synchronization invariants (position IDs, attention masks, KV-cache) across ragged tensors, a challenge that prior implementations fail to address. We formalize these invariants, prove they are preserved by our update rule, and present EqSpec, the first algorithm that guarantees output distribution equivalence by enforcing them after each verification round (Section 3.1). We analyze the cost structure of batch speculative decoding and show that alignment overhead grows superlinearly with batch size (Section 3.2); empirically, realignment consumes up to 40% of computation at batch size 8 (Section 4.3). This is an inherent cost of algorithmic correctness across ragged tensors, not an implementation inefficiency. The superlinear growth of alignment overhead represents a fundamental barrier that explains why prior methods either sacrifice validity or fail to scale.
@@
To reduce this overhead, we introduce EXSpec (Section 3.2), which expands the scheduling scope beyond a single batch. EXSpec maintains a pool of active sequences and dynamically groups those with identical lengths, eliminating realignment for homogeneous groups. This strategy preserves the scaling efficiency of standard batching while retaining per-sequence speculation benefits. At batch size 8, EXSpec achieves a 3× throughput improvement over batch size 1. Beyond batch size 8, throughput degrades as sequence length diversity reduces grouping effectiveness, forcing more frequent fallbacks to realignment. Section 4.3 examines these scaling dynamics in detail.
@@
Our experiments on SpecBench (Xia et al., 2024) yield two main results. First, Algorithmic Correctness: unlike prior approaches such as BSP and DSD, which suffer severe output corruption, our method maintains algorithmic correctness across Vicuna-7B/68M (Zheng et al., 2023), Qwen3-8B/0.6B (Yang et al., 2025), and GLM-4-9B/0.6B (GLM et al., 2024) model pairs, achieving 95% decoding-equivalence with residual divergence attributable to floating-point non-determinism in GPU inference, not synchronization failures. Second, Scalability: at batch size 8, EXSpec achieves up to a $3\times$ speedup over batch size 1, successfully combining batch parallelism with per-sequence speculation gains. Our contributions:
  • We identify that existing batch speculative decoding implementations violate output equivalence due to improper handling of the ragged tensor problem. We formalize the synchronization invariants (position IDs, attention masks, KV-cache alignment) that any valid implementation must satisfy and show precisely how prior methods fail to maintain them, producing corrupted outputs.
  • We present EqSpec, the first batch speculative decoding algorithm that guarantees algorithmic correctness through explicit synchronization, with formal proofs that the synchronization invariants are preserved.
  • We introduce EXSpec, which reduces alignment overhead through cross-batch scheduling that dynamically groups same-length sequences. On SpecBench across Vicuna-7B/68M, Qwen3-8B/0.6B, and GLM-4-9B/0.6B pairs, our methods achieve up to $3\times$ throughput at batch size 8 while maintaining algorithmic correctness.
@@ 2 Design Space Analysis
When sequences within a batch accept different numbers of draft tokens during verification, tensors become irregular, violating GPUs’ requirement for rectangular layouts—this is the ragged-tensor problem illustrated in Figure 2. Despite batching’s centrality to production deployments, existing implementations lack a principled design that preserves output equivalence with standard decoding while scaling with batch size. We identify three approaches to handle raggedness: Masking, Rollback, and Dynamic Padding. Yet, as our systematic analysis shows, current instantiations of these approaches fail to simultaneously maintain algorithmic correctness and performance at scale. To close this gap, we first analyze the pitfalls of each approach and then introduce the first valid batch speculative decoding with explicit synchronization requirements.
@@
This approach operates directly on ragged tensors by masking rejected tokens in attention and reassigning position IDs so new tokens align with their content positions. Across verification rounds with varying rejections, sequences accumulate padding in various positions (middle and right), forming non-contiguous position IDs that standard Transformer implementations handle poorly. BSP (Su et al., 2023) attempts this via masking but fails to maintain position-ID consistency across iterations, yielding corrupted outputs (Figure 1). EAGLE’s experimental batching code [^1] While EAGLE’s main contribution concerns improved draft models rather than batching, its repository includes experimental batch-related code we analyzed for implementation challenges. https://github.com/SafeAILab/EAGLE/issues/250 (Li et al., 2025) encounters similar framework limitations. BASS (Qian et al., 2024) addresses this through custom attention kernels that tolerate ragged padding in the middle and right, but this requires coupled modifications to both attention computation and KV-cache indexing, sacrificing portability across model architectures and attention variants.
@@
After each verification step, all sequences are truncated to the batch’s minimum accepted length (Wolf et al., 2020; kamilakesbi, 2024). This keeps the batch aligned, but it throws away additional verified tokens that faster sequences had already accepted. As batch size grows and acceptance-rate variance widens, the waste compounds; in the extreme, one persistently rejecting sequence forces single-token progress for the entire batch. In effect, throughput collapses to that of the worst-performing sequence, undermining speculative gains and rendering the approach impractical at larger scales. A variant approach, employed by vLLM’s v0 engine, avoids truncation through batch expansion: each sequence is duplicated $K$ times with progressively longer draft prefixes (e.g., $t_{1}],[t_{1},t_{2}],\ldots$), all verified in a single forward pass. While this preserves accepted tokens, it incurs $K\times$ redundant computation and memory overhead that scales with both batch size and draft length. At scale, this memory explosion caused GPU out-of-memory failures, ultimately leading to the deprecation of batch expansion in vLLM v1 (see Appendix C for details).
✓ Dynamic Padding Approach.
This approach realigns sequences after each verification by adjusting left padding to maintain right alignment, preserving all accepted tokens. While conceptually simple, algorithmic correctness requires tight synchronization of position IDs, attention masks, and the KV-cache. DSD’s experimental code (Yan et al., 2025) follows this idea but merely repads at each step—adding varying left padding without ever unpadding—thereby inflating sequences. It also contains three critical errors: (i) sampling bonus tokens from the draft model rather than the target model; (ii) redundantly regenerating KV-cache entries, causing memory bloat; and (iii) desynchronizing padding, position IDs, and the KV-cache across iterations. Despite the overhead of repeated realignment, a valid dynamic-padding implementation fits within standard frameworks and preserves all verified tokens.
Among the three approaches, only dynamic padding is viable. (1) Masking with non-contiguous position IDs addresses only half the problem: even if attention is patched to tolerate such layouts (via adjusted position IDs and masking), algorithmic correctness still requires that the KV-cache correspond to the same post-verification token stream across verification rounds. If one fully fixes both sides, the implementation must perform per-round cache/token realignment (i.e., compacting/realigning the batch state after verification), which effectively converges to dynamic padding. (2) Rollback preserves alignment but discards verified tokens by truncating all sequences to the batch’s minimum accepted length, and this waste grows with batch size. (3) Dynamic padding maintains algorithmic correctness by explicitly synchronizing both position IDs and KV-cache after each verification round. We provide a detailed bug taxonomy in Appendix A, distinguishing batch-independent implementation errors (e.g., wrong sampling distribution) from ragged tensor synchronization failures that only manifest at batch size >1>1. We identify and address all of these, but the synchronization problem is the core challenge: no prior work formalizes what invariants batch speculative decoding must maintain, and this formalization is our central contribution.
Figure 3: EqSpec synchronizes via unpad–append–repad.
Output Equivalence.
We formalize the validity requirement for batch speculative decoding. Under greedy decoding (temperature=0{}=0), speculative decoding must yield outputs identical to standard autoregressive generation, and any token-level divergence indicates an implementation error. For temperature>0{}>0, where outputs are sampled stochastically, speculative decoding must preserve the output distribution of the target model; that is, the probability of generating any token sequence must remain unchanged. This losslessness, whether measured by exact output equivalence or distributional equivalence, is the defining criterion of valid speculative decoding and the hallmark of an authentic implementation. Methods that achieve high throughput while violating this criterion, as we demonstrate for BSP and DSD in Figure.1, do not constitute valid speculative decoding regardless of their reported performance metrics.
3 Method
We propose a batch speculative decoding method that explicitly maintains synchronization across sequences (position IDs, attention masks, and KV-cache) while also aiming to keep the added synchronization overhead small for higher throughput. The main trade-off is that maintaining equivalence to standard decoding requires re-synchronizing position IDs, attention masks, and the KV-cache after each verification round when the batch becomes ragged, and this synchronization can account for up to 40% of total computation. We introduce two complementary mechanisms: EqSpec specifies and enforces the minimal synchronization invariants required for valid computation (Section 3.1), while EXSpec groups same-length sequences to avoid synchronization overhead (Section 3.2).
3.1 EqSpec: Algorithmic Correctness via Synchronization Invariants
Rectangular Alignment.
We formalize the algorithmic correctness requirements for batch speculative decoding. Let $B$ denote the batch size, and let sequence $i$ have padding length $p_{i}$, content length $c_{i}$, and $a_{i}$ newly accepted tokens (including the bonus token) after verification round $t$. A valid implementation must maintain rectangular alignment across the batch:
$\quad\forall i\in[1,B],\quad p_{i}^{(t)}+c_{i}^{(t)}=L^{(t)}$ (I1)
where $L^{(t)}$ is the padded sequence length at round $t$.
Position-ID Contiguity.
Additionally, position IDs must be derived consistently from the attention mask:
$\quad\mathrm{pos}_{i}[j]=\max\left(0,\sum_{k=0}^{j}\mathbf{1}[\mathrm{tok}_{i}[k]\neq\texttt{pad}]-1\right)$ (I2)
which guarantees contiguous position IDs starting from zero at the first content token, as required by RoPE (Su et al., 2024) and other position-aware attention mechanisms. Continuous batching systems (e.g., vLLM, SGLang) process requests using variable-length packing, sidestepping I1, but still require scatter-gather across requests and rollback both position-ID and KV-cache for rejected tokens—a one-dimensional analog of alignment. Critically, I2 remains a prerequisite for algorithmic correctness regardless of batching strategy; note that neither vLLM nor SGLang currently supports continuous batching speculative decoding with external draft models (see Appendix.E).
After batch verification, sequence $i$ accepts $a_{i}$ tokens, breaking rectangular alignment since $a_{i}$ varies across sequences. To restore I1, the new padding must satisfy:
$\forall i\in[1,B],\quad p_{i}^{(t+1)}+c_{i}^{(t)}+a_{i}=L^{(t+1)}$ (1)
where $L^{(t+1)}=\max_{j}(c_{j}^{(t)}+a_{j})$. This yields the per-sequence padding offset:
$\delta_{i}=p_{i}^{(t+1)}-p_{i}^{(t)}=\left(L^{(t+1)}-L^{(t)}\right)-a_{i}$ (2)
We formally prove that this update rule preserves rectangular alignment I1 across all verification rounds (Theorem B.1 in Appendix). Invariant I2 is enforced by recomputing position IDs after each repadding operation.
Algorithm 1 EqSpec: Minimalism Batch Spec
0:  Draft model $\mathcal{M}_{d}$, Target model $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Max tokens $T$, Draft length $K$
0:  Generated sequences $\mathcal{S}$
@@ 3.1 EqSpec: Algorithmic Correctness via Synchronization Invariants
1:  $\mathcal{S}\leftarrow\texttt{Tokenize}(\mathcal{P})$ $\triangleright$ Batch left padding
2:  $\textit{KVCache}\leftarrow\emptyset$
3:  while until max new tokens do
4:  Phase 1: Draft Generation
5:  $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{S},K)$
6:  Phase 2: Batch Verification
7:  $\triangleright$ See Alg. 3 in Appendix for BatchVerify
8:  $A,B,\textit{KVCache}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{S},D,\textit{KVCache})$
9:  $\triangleright$ See Figure 3 for illustration on index offset.
10:  Phase 3: Unpad-Append-Repad
11:  for each sequence $i$ in batch do
12:   $\mathcal{S}[i]\leftarrow\texttt{Unpad}(\mathcal{S}[i])$
13:   $\mathcal{S}[i]\leftarrow\mathcal{S}[i]\oplus A[i]\oplus B[i]$
14:  end for
@@
17:  end while
18:  return $\mathcal{S}$
Unpad-Append-Repad.
Figure 3 illustrates how EqSpec restores a valid batch after each verification round. The procedure applies the offset $\delta_{i}$ to realign all batch state: token sequences receive fresh left padding, attention masks are recomputed, position IDs are recalculated via I2, and KV-cache entries are shifted by $\delta_{i}$ positions to maintain token-to-cache correspondence.
During batch verification, the target model performs a single forward pass over draft tokens, detects the first mismatch per sequence, and samples a bonus token from the target distribution (Algorithm 3 in Appendix.C). However, since this sampling occurs after the forward pass completes, the bonus token lacks KV-cache entries and must be included in the next forward pass to create them. The complete EqSpec procedure is detailed in Algorithm 1.
@@
Algorithm 2 EXSpec: Cross-Batch Scheduling
0:  Draft and Target model $\mathcal{M}_{d}$, $\mathcal{M}_{t}$, Prompts $\mathcal{P}$, Window size $W$, Batch size $B$
0:  Generated sequences $\mathcal{S}$
1:  $\textit{Pool}\leftarrow\texttt{InitSequencePool}(\mathcal{P})$
2:  $\triangleright$ Tokenize and optionally sort by length
3:  $\textit{Window}\leftarrow\texttt{RefillWindow}(\textit{Pool},W)$
4:  while Pool.hasActive() do
5:  Phase 1: Lazy Realignment
6:  $\mathcal{B},\textit{mask},\textit{KV}\leftarrow\texttt{GetBatch}(\textit{Window},B)$
7:  $\triangleright$ Try same-length concatenation
8:  $\triangleright$ Fallback to Unpad-Repad Realignment
9:  Phase 2: Draft Generation
10:  $D\leftarrow\mathcal{M}_{d}.\texttt{Generate}(\mathcal{B},\textit{mask},K)$
11:  Phase 3: Batch Verification
12:  $\triangleright$ See Alg. 3 in Appendix for BatchVerify
13:  $A,B,\textit{KV}\leftarrow\texttt{BatchVerify}(\mathcal{M}_{t},\mathcal{B},D,\textit{KV})$
14:  Phase 4: Write-Back and Window Refill
15:  for $i\in\mathcal{B}$ do
16:   $\textit{Pool}[i]\leftarrow\textit{Pool}[i]\oplus A[i]\oplus B[i]$
17:   $\textit{Pool.KV}[i]\leftarrow\textit{KV}[i]$
18:   if $\textit{isComplete}(\textit{Pool}[i])$ then $\textit{Pool.deactivate}(i)$
19:  end for
21:  end while
22:  return Pool.sequences
3.2 EXSpec: Reducing Overhead via Cross-Batch Scheduling
Theoretical Speedup Analysis
The speedup of speculative decoding depends on both the token-acceptance rate and computational costs. The original formulation by Leviathan et al. (2023) analyzes single-sequence performance, modeling the expected tokens generated per iteration as $(1-\alpha^{k+1})/(1-\alpha)$, where $\alpha$ is the token acceptance rate (TAR) and $k$ is the number of draft tokens per speculation round. This single-sequence view assumes no batch overhead and focuses purely on acceptance dynamics. For batch speculative decoding, we introduce a batch-aware speedup formula: